Steve has deposited $50,000 into an account earning 6% compounded quarterly. After 5 years, he transfers the money in his account into a RRIF earning 4.5% compounded monthly, and makes beginning-of the month withdrawals of $2000. How long can Steve make these withdrawals?
1
Expert's answer
2020-03-31T08:51:06-0400
  1. Find the accumulated amount

"FV=PV\times(1+\frac{r}{n})^{nm}=50000\times(1+\frac{0.06}{4})^{20}=\\=50000\times(1+0.015)^{20}=50 000 \times1.346855=67342.75"

2. Find the term:

Since Steve transfers money from his account to RRIF, this turns out to be annuity:

Therefore, we apply the formula:

"S=R \times \frac{(1+i)^n-1}{i}"

Therefore, we apply the formula

"\frac{S}{R}\times i+1=(1+i)^n"

Let us take the natural logarithm of this equality

"n\times \ln(1+i)=\ln(\frac{S}{R}\times i+1)"


"n=\frac{\ln(\frac{S}{R}\times i+1)}{\ln(1+i)}"


"n=\frac{\ln(\frac{67342.75}{2000}\times0.00375+1)}{\ln(1+0.00375)}"


"n=\frac{\ln(33.671375\times0.00375+1)}{\ln1.00375}"


"n=\frac{0.1188}{0.00374}=31.76"


32 months